OpenStudy (anonymous):

there were 24 people, teachers and students, in the Christmas play. there were three times as many students as teachers. which system of equations could be used to find the number of teachers,t, and the number of students,s, in the play?

OpenStudy (anonymous):

its writing systems of equations

OpenStudy (aaronandyson):

what are you're options?

OpenStudy (anonymous):

there is none

OpenStudy (anonymous):

is this your first time on openstudy?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

welcome to openstudy then :)

OpenStudy (aaronandyson):

Are you given any options to the question? @chloelynn_

OpenStudy (anonymous):

no, its just a solve by yourself question

OpenStudy (anonymous):

you can do this without a system of equations

OpenStudy (anonymous):

you can do it with one too if you like say S is students, T is teachers, then since the total is 24 you can say \[S+T=24\]

OpenStudy (anonymous):

then three times as many students as teachers means \[3T=S\]

OpenStudy (anonymous):

so u do matrix?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[S+T=24\\ 3T=S\] plug the scone in to the first and get \[S+3S=24\] or \[4S=24\]